In the determination of Young’s modulus Y=4MLg/pi*ld^2 by using Searle’s method, a wire of length L = 2mand diameter d = 0.5 mm is used. For a load M = 2.5 kg, an extension l = 0.25 mm in the length of the wireis observed. Quantities d and l are measured using a screw gauge and a micrometer, respectively. They have the same pitch of 0.5 mm. The number of divisions on their circular scale is 100. The contributions to the maximum probable error of the Y measurement

Grade:9

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,

L.C. = 0.5/100= 0.005 mm
∆Y/y = ∆ll + 2∆(d)/d
∆l/l=0.005×10^{-3}/0.25×10{-3}=1/50
2∆d/d=2*0.005×10^{-3}/0.5×10{-3}=1/50
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty